### v1

# A Lightweight Method for Defending Against UAF Vulnerabilities

Abstract.
The widespread presence of Use-After-Free (UAF) vulnerabilities poses a serious threat to software security, with dangling pointers being considered the primary cause of these vulnerabilities. However, existing methods for defending against UAF vulnerabilities by eliminating dangling pointers need to interrupt the program’s execution when encountering pointer assignment operations to look up the objects pointed to by the pointers and store the memory addresses of the pointers in a specific data structure. This makes these methods not lightweight. To overcome this drawback, we propose a novel approach called LightDE. This method does not require storing the memory addresses of pointers or locating the objects pointed to by pointers during program execution. LightDE uses our proposed structure-sensitive pointer analysis method to determine the objects pointed to by pointers and stores the pointing relationships in the program’s data segment during program compilation. Since LightDE only needs to check whether the pointers identified by the pointer analysis point to the released objects when the objects are released, LightDE is very lightweight. Our experimental results show that LightDE can effectively defend against UAF vulnerabilities, and the additional performance overhead it introduces is very low.
Use-after-free, Dangling pointer, Pointer analysis, LLVM
copyright: nonecopyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06ccs: Security and privacy Software security engineering
1. Introduction
Use-After-Free (UAF) vulnerabilities pose a serious threat to program security as they can lead to arbitrary code execution, privilege escalation, and data leakage (Gui et al., 2021; Song et al., 2019; Szekeres et al., 2013). Unfortunately, the number of UAF vulnerabilities has been increasing over the years, and their impact has become increasingly severe (Gui et al., 2021; Yan et al., 2018; Cuoq et al., 2012; Ye et al., 2014). Table 1 shows the changes in the number of UAF vulnerabilities in the NVD database from 2011 to 2023, as well as the number of high-threat UAF vulnerabilities. There are two types of UAF vulnerabilities: stack UAF and heap UAF (Younan, 2015; Van Der Kouwe et al., 2018; Szekeres et al., 2013). Stack UAF occurs when a dangling pointer is used to access stack memory that has already been freed. However, modern compilers generate warning messages for such vulnerabilities, making it easy for programmers to detect and fix them. Additionally, stack space allocation and deallocation are automatic, making stack UAF vulnerabilities difficult to exploit. Heap UAF refers to accessing freed heap memory through a dangling pointer. The difficulty in defending against heap UAF primarily stems from the flexible allocation and deallocation of heap memory, as well as the complexity of pointer propagation. All UAF vulnerabilities discussed in this paper are heap UAF.
Figure 1. Use-after-free Vulnerabilities in NVD (NIST, 2005).
Researchers have proposed various methods to defend against UAF vulnerabilities, but each approach has its limitations and drawbacks. One method involves checking the validity of pointer access during pointer dereferencing, which requires verifying memory access legality before each pointer dereference (Nagarakatte et al., 2010, 2012; Necula et al., 2002; Condit et al., 2003; Zhang et al., 2019). However, this method can lead to incompatibility with existing libraries and introduce unacceptable runtime overhead. An alternative approach modifies memory allocation and deallocation strategies, making it harder for attackers to exploit UAF vulnerabilities (Boehm et al., 2002; Oiwa, 2009; Zhang and Zonouz, 2018; Novark and Berger, 2010; Berger and Zorn, 2006; Dang et al., 2017). However, this method only provides probabilistic defense, and attackers can bypass it through techniques like heap spraying and other attack methods. Recent research has proposed defending against UAF vulnerabilities by eliminating dangling pointers (Lee et al., 2015; Younan, 2015; Van Der Kouwe et al., 2017; Liu et al., 2018; Shen and Dolan-Gavitt, 2020). Eliminating dangling pointers is a very promising method for mitigating UAF vulnerabilities, as dangling pointers are the root cause of these vulnerabilities. However, existing methods of eliminating dangling pointers need to record pointer memory addresses and look up the objects pointed to by pointers during program execution, making these methods not lightweight.
Since dangling pointers are the root cause of UAF vulnerabilities (Lee et al., 2015; Younan, 2015; Van Der Kouwe et al., 2017; Liu et al., 2018; Shen and Dolan-Gavitt, 2020), this paper focuses on defending against UAF vulnerabilities by eliminating dangling pointers. Current methods (Lee et al., 2015; Younan, 2015; Van Der Kouwe et al., 2017; Liu et al., 2018; Shen and Dolan-Gavitt, 2020) for defending against UAF vulnerabilities by eliminating dangling pointers, rely on inserting functions after pointer assignment instructions during compilation to track pointers and the objects they reference. This method requires the program to pause its normal execution flow when encountering a pointer assignment instruction and to execute the logic of the inserted function. The role of the inserted code is to look up the objects pointed to by the pointers and record the memory addresses of the pointers in the object’s pointer set. Although existing dangling pointer elimination schemes use different methods to locate the objects pointed to by pointers and various data structures to store the memory addresses of pointers to improve program execution efficiency, the basic principle remains the same. In this process, locating the objects pointed to by pointers and recording the memory addresses of pointers are time-consuming, making current dangling pointer elimination methods not lightweight.
To address the issue of existing methods being unable to eliminate dangling pointers in a lightweight manner, we propose LightDE, a lightweight approach to eliminate dangling pointers and defend against UAF vulnerabilities. LightDE’s uniqueness lies in its combination of pointer analysis (Lhotak et al., 2013; Hardekopf and Lin, 2007; Pereira and Berlin, 2009; Kroening and Tautschnig, 2014) to determine the objects pointed to by pointers at compile time, and storing the pointing relationships in the program’s data segment during compilation. The significant advantage of our method over existing methods is that it does not require locating the objects pointed to by pointers or storing the memory addresses of pointers during program execution. Additionally, we use an asynchronous method to eliminate dangling pointers by running a separate thread on an idle CPU core. This thread uses the pointing relationships stored in the program’s data segment to check if the pointers identified by the pointer analysis point to the released objects and eliminate the corresponding dangling pointers.
We implemented a prototype system of LightDE and demonstrated its effectiveness using real UAF vulnerabilities. Our experimental results on the SPEC CPU2006 (Henning, 2006) benchmark suite show that LightDE introduces very low additional runtime overhead, with a geometric mean of only 8.66%. The geometric mean of the additional memory overhead introduced by LightDE is 7.26%. Experimental results on the SPLASH-2X (Group et al., 2011) benchmark suite show that LightDE scales well on multithreaded applications.
To summarize, our contributions in the field of defending against UAF vulnerabilities are as follows:
1. We propose a lightweight method for eliminating dangling pointers, LightDE. Our approach determines pointer relationships during program compilation and embeds this information into the program’s data segment at compile time. This allows our method to avoid interrupting the program’s execution flow to locate pointer targets or store the memory addresses of pointers, making LightDE very lightweight.
2. We propose a structure-sensitive pointer analysis method to obtain the structural information of objects and improve the accuracy of pointer analysis results.
3. We developed a prototype system of LightDE based on LLVM 13 and SVF-2.5 to validate our design. We conducted an experiment to evaluate the LightDE prototype. The experimental results show that LightDE can effectively defend against UAF vulnerabilities while introducing very low additional performance overhead.
2. Motivation Example
1typedef struct {
2 int* arr;
3} A;
4void main() {
5 A* a = (A*)malloc(8); //o1
6 storePtr(&a);
7 int* b = (int*)malloc(4); //o2
8 storePtr(&b);
9 a->arr = b;
10 storePtr(&a–>arr);
11
12 free(b);
13 = *a->arr;
14}
15
16void storePtr(void** ptrAddr) {
17 // look up the pointer pointing to object
18 // store the pointer address to the global data structure
19
20}
Figure 2. A Motivation Example.
Figure 3. Overview of LightDE.
To further illustrate the issue that existing methods of eliminating dangling pointers (Lee et al., 2015; Younan, 2015; Van Der Kouwe et al., 2017; Liu et al., 2018; Shen and Dolan-Gavitt, 2020) are not lightweight, we use the example in Figure 2 to explain this problem. In this example, a type $A$ object $o1$ is created on line 5, and the memory address of $o1$ is stored in the variable $A*\ a$. On line 7, an integer object $o2$ is allocated on the heap, and the memory address of $o2$ is stored in the variable $int*\ b$. On line 9, the memory address of $o2$ is assigned to the first field of object $o1$. The code segment highlighted in red in Figure 2 is an abstraction of the inserted instructions in existing dangling pointer elimination methods. The function $storePtr$ in Figure 2 also abstracts the insertion function from existing dangling pointer elimination methods. The primary responsibility of this function is to locate the object pointed to by a pointer and record the pointer’s address in the designated global data structure. It is worth noting that various methods may have slight differences in the specific operations of this function, but the basic functionality is similar. That is, tracking the pointers and the objects they point to in order to eliminate the corresponding dangling pointers when memory is freed. When the program executes the code in line 6 of Figure 2, the $storePtr$ function locates the object pointed to by $A*\ a$ and records the memory address of $A*\ a$ in the pointer set of the corresponding object. In this example, the pointer $A*\ a$ will be recorded in the pointer set of the object $o1$. Similarly, the code on lines 8 and 10 will trigger similar processes. This design means that whenever the program encounters a pointer assignment operation, it will perform object lookup and pointer memory address recording operations. This severely impacts the program’s performance. Through the example in Figure 2, it is easy to understand why existing dangling pointer elimination methods are not lightweight.
3. Overview of LightDE
We illustrate the working process of LightDE in Figure 3. First, we compile the program’s source code into LLVM Intermediate Representation (LLVM IR) using the LLVM compiler. We then perform pointer analysis based on the LLVM IR. Our structure-sensitive pointer analysis (Balatsouras and Smaragdakis, 2016; Barbar et al., 2020; Pearce et al., 2007; Avots et al., 2005) is conducted in two phases. The primary purpose of the first phase of pointer analysis is to identify the objects pointed to by pointers and the structural information of these objects. The second phase of the pointer analysis incorporates the structural information of objects obtained in the first phase and uses this structural information to filter out spurious pointers pointing to object fields. Next, based on the results of the pointer analysis, we categorize pointers into three types: heap pointers, global pointers, and stack pointers. This classification is essential because the elimination method for each type of pointer is different. Based on the results of the pointer analysis, we embed the pointer points-to information in the program’s data segment. However, it is not possible to determine the runtime memory address of a pointer in the static case. We will explain our method for storing pointer points-to relationships at compile time and mapping the results of static pointer analysis to runtime information in Section 4.2.
In the pointer analysis phase, our method of abstracting objects is to treat each heap allocation point as an object. This abstraction method is too coarse for eliminating dangling pointers, so we need to accurately identify each object allocated and released by the program. To address this issue, we use a multithreaded hash table (Gao et al., 2005) to map static objects to runtime objects. We insert a function after each object allocation point in the program to store the allocated object information and the static information of the object in our multithreaded hash table. Similarly, we insert a function before each object release point to identify the released objects and pass their information to the LightDE thread. We remove the program’s original memory release function, allowing the LightDE thread to handle memory space release to prevent UAF vulnerabilities during the elimination of dangling pointers by the LightDE thread. The creation of the LightDE thread occurs when the program starts execution. The function of this thread is to eliminate the corresponding dangling pointers based on the pointing relationships stored in the program’s data segment after receiving information about released objects passed from the program’s original thread. Finally, the LLVM IR with the inserted data and functions is compiled into a binary program.
4. Design
In this section, we first explain LightDE’s structure-sensitive pointer analysis method to obtain the pointing relationships of pointers and the structural information of objects. Then, we explain how LightDE maps the pointer analysis results to runtime information to accurately eliminate dangling pointers.
4.1. Structure-sensitive Pointer Analysis
In this section, we first explain our method for abstracting programs, then describe our pointer analysis method.
4.1.1. Program Abstraction
This section includes three parts: abstraction of program domains, abstraction of program instructions, and symbols used in pointer analysis.
Table 1. Analysis Domains of Our Pointer Analysis.
Analysis domains
$t$in $\mathcal{T}$ program types
$f$in $\mathcal{C}$ field offsets
$fld$in $\mathcal{F}$ fields
$c$in $\mathcal{C}$ constants
$s$in $\mathcal{S}$ virtual registers
$g$in $\mathcal{G}$ global variables
$m$in $\mathcal{M}$subseteq $\mathcal{G}$ program functions
$p,q,x,a$in $\mathcal{P}$ $=$ $\mathcal{S}$cup $\mathcal{G}$ top-level variables
$o$in $\mathcal{O}$ abstract objects
[…]
$d$in $\mathcal{A}$ $=$ $\mathcal{O}$cup $\mathcal{Y}$ address-taken variables
$v$in $\mathcal{V}$ $=$ $\mathcal{P}$cup $\mathcal{A}$ program variables
The domains related to our pointer analysis are shown in Table 1, mainly including variables, functions, types, and objects. Variables in the program $\mathcal{V}$ can be divided into two types: address-taken variables $\mathcal{A}$, which can be accessed through pointers, and top-level variables $\mathcal{P}$, which are never accessed through pointers. Address-taken variables $\mathcal{A}$ include abstract objects $\mathcal{O}$ and fields of abstract objects $\mathcal{Y}$. Heap objects are a type of address-taken object, and our method for abstracting heap objects is to treat each heap allocation point as a distinct object. We represent object fields by creating a sub-object for each field and establishing a relationship between the sub-object and its parent object. The creation of sub-objects is based on the offset of their fields relative to the base object.
Table 2. LLVM-like Instructions.
LLVM-like instructions
$p=alloca\ t,nbytes$ STACK/GLOBAL
$p=malloc\ nbytes$ HEAP
$p=q$ COPY
$p=(t)q$ CAST
$p=*q$ LOAD
$*p=q$ STORE
$p=\phi(q,h)$ PHI
$p=\&q\rightarrow fld$ FIELD
$p=q(...,a,...)$ CALL
$ret\ p$ RET
Our pointer analysis is based on the LLVM IR of programs. LLVM defines a rich set of instructions, but only a small subset is relevant to our pointer analysis. Table 2 shows the set of instructions related to our pointer analysis. STACK/GLOBAL instructions allocate stack or global objects. HEAP instructions are used for heap object allocation. The COPY instruction is an assignment statement. PHI is the convergence point of a branching structure that selects different values based on the branch taken. CAST is used to interpret an object as a specific type to satisfy LLVM’s type checking without altering the contents in memory. LOAD and STORE are operations for reading from or writing to address-taken objects. FIELD obtains the address of an object’s field. CALL and RET instructions are used for function calls and returns.
The symbols used in this paper for pointer analysis are as follows: $pt(p)$ represents the set of objects pointed to by pointer $p$. For example, $pt(p)=\{o1,o2\}$ means that pointer $p$ points to objects $o1$ and $o2$. $pt.t(o)$ represents the set of types owned by object $o$. $pt.o(t)$ represents the set of field offsets that type $t$ owns. The function $off()$ is used to obtain the offset of a field within an object. The symbol $\widehat{o_{f}}$ denotes the sub-object of object $o$ with an offset of $f$. The symbol * represents any field offset. The rest of the symbols have the same meaning as mathematical set operators.
4.1.2. Pointer Analysis
The instructions involved in pointer analysis can be categorized into five types: object creation instructions, object propagation instructions, type conversion instructions, field access instructions, and function call and return instructions. The inference rules for the first phase of pointer analysis are shown in Figure 4.
$\begin{array}[]{l}\textbf{{\scriptsize[STACK/GLOBAL]}}\\ $$\ \ p=alloca\ t,nbytes$$\\ \hline\cr$$\ \ pt(p)=pt(p)\cup\{o\}\\ pt.t(o)=pt.t(o)\cup\{t\}$$\end{array}$ $\quad\ \ \ \ \quad\ \ \quad\ \ \begin{array}[]{l}\textbf{{\scriptsize[HEAP]}}\\ $$\ \ p=malloc\ nbytes$$\\ \hline\cr pt(p)=pt(p)\cup\{o\}$$\end{array}$ $\quad\ \ \quad\ \ \quad\ \ \begin{array}[]{l}\textbf{{\scriptsize[COPY]}}\\ $$\ \ \ \ \ \ p=q$$\\ \hline\cr$$pt(q)\subseteq pt(p)$$\end{array}$ $\quad\ \ \ \ \quad\ \ \quad\ \ \begin{array}[]{l}\textbf{{\scriptsize[CAST]}}\\ $$p=(t)q\ \ \ \ \ o\in pt(q)$$\\ \hline\cr$$\ \ \ \ \ pt(q)\subseteq pt(p)\\ pt.t(o)=pt.t(o)\cup\{t\}$$\end{array}$
$\begin{array}[]{l}\textbf{{\scriptsize[PHI]}}\\ $$\ \ \ \ \ \ \ \ p=\phi(q,x)$$\\ \hline\cr$$pt(q)\cup pt(x)\subseteq pt(p)$$\end{array}$ $\quad\ \ \quad\ \ \quad\ \ \begin{array}[]{l}\textbf{{\scriptsize[LOAD]}}\\ $$p=*q\ \ o\in pt(q)$$\\ \hline\cr$$\ \ pt(o)\subseteq pt(p)$$\end{array}$ $\quad\ \ \quad\ \ \begin{array}[]{l}\textbf{{\scriptsize[STORE]}}\\ $$*p=q\ \ o\in pt(p)$$\\ \hline\cr$$\ \ pt(q)\subseteq pt(o)$$\end{array}$ $\quad\ \ \quad\ \ \quad\ \ \quad\ \ \begin{array}[]{l}\textbf{{\scriptsize[FIELD-VAR]}}\\ $$p=\&q\rightarrow*\ \ \ o\in pt(q)$$\\ \hline\cr$$\ \ pt(p)=pt(p)\cup\{\widehat{o_{\ast}}\}$$\end{array}$
$\begin{array}[]{l}\textbf{{\scriptsize[FIELD]}}\\ $$p=\&q\rightarrow fld\ \ \ f=off(q\rightarrow fld)\ \ \ o\in pt(q)$$\\ \hline\cr$$\ \ \ \ \ \ \ \ \ \ \ \ \ \ pt(p)=pt(p)\cup\{\widehat{o_{f}}\}$$\end{array}$ $\quad\ \ \begin{array}[]{l}\textbf{{\scriptsize[CALL]}}\\ $$p=q(...,a,...)\ \ m(...,a^{\prime}...)\ \ m\in pt(q)$$\\ \hline\cr$$\ \ \ \ \ \ \ \ \ \ \ \ pt(a)\subseteq pt(a^{\prime})$$\\ \end{array}$ $\begin{array}[]{l}\textbf{{\scriptsize[RET]}}\\ $$p=q(...)\ \ \ ret\ p^{\prime}\ \ \ m\in pt(q)$$\\ \hline\cr$$\ \ \ \ \ \ \ pt(p^{\prime})\subseteq pt(p)$$\end{array}$
Figure 4. Stage 1 Pointer Analysis Rules.
Object creation involves two rules: [STACK/GLOBAL] and [HEAP]. [STACK/GLOBAL] is used to create stack and global objects, and it can obtain the type information of objects from the creation instruction. In this rule, an object $o$ is created, and we use $o\in pt(p)$ to indicate that pointer $p$ points to object $o$. The type of object $o$ in this rule is $t$, and we use $t\in pt.t(o)$ to indicate that the object has type $t$. The [HEAP] instruction only involves the creation of objects.
Object propagation involves the four rules [COPY], [PHI], [LOAD], and [STORE]. These rules only involve the propagation of objects. Both the [COPY] and [PHI] rules are simple assignment instructions, propagating the points-to information of the right-hand pointer to the left side. Specifically, the [PHI] rule merges the points-to sets of the two pointers on the right and propagates them to the left. [LOAD] and [STORE] propagate objects with the help of a third pointer.
Type conversion only involves the [CAST] rule. Here, we briefly explain our type model. To ensure the accuracy and completeness of object type information, an object’s type set is only allowed to grow and not shrink. We adopt this method because various types of pointers pointing to an object do not disappear when the object’s type changes. These pointers can still access fields of the object’s previous types. Thus, we cannot discard an object’s existing types when its type changes. In the [CAST] rule, the target type is added to the object’s type set.
Field access instructions involve the [FIELD] and [FIELD-VAR] rules. [FIELD] is used to analyze cases where an object’s field is accessed by a known offset. In such cases, the pointer points to the offset of the corresponding field of the object. [FIELD-VAR] is used to analyze cases where an object’s field is accessed by an unknown offset. In these cases, we conservatively assume that the pointer might point to any field of the object. Additionally, it is worth noting that we treat arrays as a special kind of structure. The elements within an array have the same type, but each element has a different offset. Therefore, our method is also array-sensitive.
Function calls involve the [CALL] and [RET] rules. We employs an On-The-Fly approach for resolving function calls. The On-The-Fly further analyzes the call relationships between functions based on the results of pointer analysis. This kind of function call analysis can effectively analyze indirect function calls, such as $call$ $*q$.
$\begin{array}[]{l}\textbf{{\scriptsize[FIELD]}}\\ $$\ \ \ \ \ \ \ \ p=\&q\rightarrow fld\ \ \ \ \ \ \ \ \ \ f=off(q\rightarrow fld)$$\\ \hline\cr$$\forall o_{i}\in pt(q)\land t\in pt.t(o_{i})\land f\in pt.o(t),pt(p)=pt(p)\cup\{\widehat{o_{f}}\}$$\\ \end{array}$
Figure 5. Stage 1 Pointer Analysis Rules.
The second phase of pointer analysis uses the object type information obtained in the first phase to filter out spurious pointers pointing to object fields. The inference rules in the second phase are similar to those in the first phase, with only the [CAST] and [FIELD] rules being different. Since the type information of objects has already been obtained in the first phase, the [CAST] rule is not needed in the second phase. Type conversion instructions are treated as copy instructions. The [FIELD] rule in the second phase checks whether each type in the object’s type set contains the field offset being accessed by the current statement. If the field offset cannot be found, the points-to relationship is filtered out. If the field offset exists, a sub-object is created for that offset, and the pointer points to the created sub-object. The [FIELD] rule is shown in Figure 5.
The specific pointer analysis process involves the following steps: During stage 1 of pointer analysis, we repeatedly apply the pointer analysis rules in Figure xxx until the pointer pointing set no longer changes, marking the end of stage 1 pointer analysis. Then, stage 2 of pointer analysis begins, which is similar to stage 1 but uses the object type information obtained from the first stage. The pointer analysis concludes when the set of pointer pointing set no longer changes.
We conducted a soundness analysis of our method. When analyzing instructions that access object fields, our method uses the type information of the objects to check whether the object has a field at the corresponding offset, making the pointer analysis results more accurate. However, there may be cases where the entire object is accessed as raw memory without using structural information. For objects that do not contain structural information, we do not use structural information to filter their points-to relationships. Our method is capable of accurately analyzing both structured and unstructured objects, thereby ensuring its soundness.
4.2. Mapping Static Results to Runtime
In this section, we explain the method of mapping static analysis results to runtime information of the program. First, we describe the method of mapping static objects to runtime objects. Then, we explain the method of mapping static pointers to runtime pointers. Finally, we explain our method for accurately eliminating dangling pointers by utilizing the pointer analysis results and the program’s runtime information.
4.2.1. Object Mapping
There are two reasons why we need to map static objects to runtime objects: 1. A single static object corresponds to multiple runtime objects. 2. The memory addresses of objects can only be determined at runtime. During pointer analysis, we abstract each heap allocation point as a specific object. This abstraction method allows us to quickly analyze points-to relationships during pointer analysis, but it is not precise enough for eliminating dangling pointers. This is because the same allocation point can create different objects in different contexts. Therefore, it is unreasonable to continue using this abstraction method when eliminating dangling pointers.
To distinguish different objects created at the same allocation point, we use the runtime information of the program. We insert a function after each allocation point in the program. The main function of this inserted function is to store the allocated memory information in our designed multithreaded hash table and to pass the information of the created objects to our thread through a multithreaded lock-free queue (Aldinucci et al., 2012). The LightDE thread groups all dynamic objects created at the same allocation point together, enabling us to find the corresponding runtime objects based on the static object. The object information stored in our data structure includes the memory start address, end address, size, and the identifier of the object’s corresponding static object.
4.2.2. Pointer Mapping
In this section, we explain our method for mapping static pointers to runtime pointers.
Heap Pointer Mapping Before introducing our method for mapping heap pointers, we use an example to illustrate the relationship between static heap pointers and runtime heap pointers. Suppose there are two objects, $o1$ and $o2$. The $f$ field of $o2$ points to the object $o1$. In static analysis, $o1$ and $o2$ correspond to two allocation points, which may correspond to many runtime objects. Our mapping method involves allocating a data structure for each static object based on the pointer analysis results during program compilation. The data structure records the heap pointers pointing to the object. Since the addresses of heap pointers cannot be determined during compilation, we record the static identifier of the object containing the heap pointer and the offset of the heap pointer within the static object in the data structure. When an object is released, our method for eliminating heap pointers that point to the released object is to check the heap pointers at the same offset location in the runtime objects corresponding to the static object containing the heap pointer. If we cannot determine the exact offset of the heap pointer within the object, we check all the object’s pointer fields to see if they point to the freed object. Since our pointer analysis is structure-sensitive, we can determine all the offsets of the object’s pointer fields.
Global Pointer Mapping In this paragraph, we explain our method for mapping static global pointers to runtime global pointers. Global pointers can be accessed from any location in the program, but their memory addresses cannot be determined during program compilation. To address this issue, we assign indices based on the names of global pointers. During compilation, we store the indices of global pointers that point to a certain static object in its pointer set. We insert a function at the beginning of the program’s startup function, which stores the memory addresses of global pointers into a global array according to their indices. Using this global array, we can check whether the global pointers identified by pointer analysis point to released runtime objects based on the objects’ static identifiers.
Stack Pointer Mapping In this paragraph, we explain our method for mapping static stack pointers to runtime stack pointers. Our method involves assigning a static identifier to each function in the program during compilation, and assigning a static identifier to all stack pointers within each function in the order they are defined. During compilation, we store the function identifier of the stack pointer, as well as the stack pointer’s static identifier within the function, into the object’s pointer set. In each function, we insert instructions to store the stack pointer into our data structure based on the static identifier within the function. When an object is released, we check whether the stack pointers identified by the pointer analysis point to the released object. Since we use a separate thread to asynchronously eliminate stack pointers, we need to avoid the stack space being freed while eliminating stack pointers. We designed a dedicated stack for this purpose. When the program needs to allocate stack pointers, it allocates space on this dedicated stack. However, the original thread of the program does not release the allocated space. When a function exits, the program’s original thread generates exit information and passes it to our thread. Our thread then releases the corresponding stack space.
4.2.3. Dangling Pointer Elimination
In this paragraph, we explain the method LightDE uses to eliminate corresponding dangling pointers after an object is released. We insert our memory management function before the release points identified by the pointer analysis and remove the program’s original memory release function. The LightDE thread releases the program’s allocated memory to prevent UAF vulnerabilities from occurring during the elimination of dangling pointers by LightDE. Our memory management function removes the information of the released object from our multithreaded hash table and passes this information to the LightDE thread. The LightDE thread then retrieves the static identifier of this runtime object and uses it to obtain the static pointers stored in the program’s data segment that point to the object. Next, the LightDE thread uses the pointer mapping method described in 4.2.2 to check whether the corresponding runtime pointers point to the released object, and it invalidates the corresponding dangling pointers.
5. Evaluation
Our experiments mainly evaluate three aspects of LightDE: effectiveness, performance overhead, and scalability in multithreaded applications. First, we experimentally assess LightDE’s effectiveness in defending against real UAF vulnerabilities. Next, we evaluate the additional performance overhead introduced by LightDE using the SPEC CPU2006 benchmark suite (Henning, 2006). We chose the SPEC CPU2006 benchmark suite because existing dangling pointer elimination methods use this suite, making it convenient for us to compare our method with existing methods. Finally, we assess LightDE’s scalability in multithreaded applications using the SPLASH-2X benchmark suite (Group et al., 2011). We believe that heap pointers and global pointers exist for a longer duration and have a wider scope of influence, making them more likely to cause UAF vulnerabilities. Therefore, in our experiments, we not only analyzed the impact on program performance when considering all pointers, including stack, heap, and global pointers, but also separately analyzed the impact on program performance when considering only heap and global pointers. We implemented a prototype system of LightDE based on LLVM-13 (llvm, 2000) and SVF-2.5 (Sui and Xue, 2016; SVF, 2022). All experiments were conducted on a machine running 64-bit Ubuntu 20.04 LTS, equipped with a 13th Gen Intel(R) Core(TM) i7-13700H processor at 2.640GHz and 32GB of RAM.
5.1. Effectiveness
Table 3. Real-world UAF Vulnerabilities Used for Evaluation.
CVE ID Application Protected
CVE-2024-24260 (CVE-2024-24260, 2024) media-server v1.0.0
CVE-2020-16590 (CVE-2020-16590, 2020) Binutils 2.35
CVE-2018-5747 (CVE-2018-5747, 2018) lrzip 0.631
CVE-2017-12448 (CVE-2017-12448, 2017) Binutils 2.29
To evaluate the effectiveness of LightDE in mitigating UAF vulnerabilities, we applied LightDE to four programs with known UAF vulnerabilities and ran exploitation programs to determine the effectiveness of LightDE in defending against these vulnerabilities. The experimental results show that LightDE was able to prevent the exploitation of UAF vulnerabilities in all four applications, as shown in Table 3.
5.2. Performance on SPEC CPU2006
In this section, we evaluate the performance overhead of LightDE on the SPEC CPU2006 benchmark suite.
5.2.1. Statistics
Table 4. Statistics for SPEC CPU2006.
Benchmark
static
object #
free
site #
static heap
pointer #
static global
pointer #
static stack
pointer #
runtime
alloc obj #
runtime
free obj #
invalidate
pointers #
400.perlbench 22 253 357 3626 2033 358M 356M 9569K
401.bzip2 6 7 18 3 0 162 144 18
403.gcc 1724 1201 2021962 281490 759291 21M 21M 2938K
429.mcf 4 9 27 10 0 3 3 4
433.milc 61 30 416 37 18 6509 6468 0
444.namd 44 402 21 6 41 1329 1321 56
445.gobmk 51 9 298 65 238 656K 656K 209K
447.dealII 388 1502 28384 12347 49348 24M 24M 14M
450.soplex 269 469 52972 13609 3145 67K 67K 473
453.povray 532 583 284014 45194 302064 2460K 2460K 1747K
456.hmmer 224 330 11754 1 538 1351K 1351K 709K
458.sjeng 10 16 27 7 0 4 0 0
462.libquantum 7 7 0 2 5 95 95 0
464.h264ref 293 582 7446 429 4 176K 176K 828K
470.lbm 2 2 0 4 0 2 2 1
471.omnetpp 2434 2155 831479 204579 49198 312M 266M 451K
473.astar 64 81 138 47 17 4790K 4790K 3049K
482.sphinx3 162 93 5225 100 9 7918K 7918K 4115K
483.xalancbmk 564 5605 253507 192478 443941 126M 126M 476K
Table 4 shows the static data of the SPEC CPU2006 benchmark suite. The second column displays the number of static objects for each project. The seventh column shows the number of objects allocated at runtime. By comparing these two columns, it is evident that a single static object corresponds to multiple runtime objects. The third column shows the number of release points identified by the pointer analysis, while the eighth column shows the number of objects released at runtime. From these two columns, it is also clear that a single release point releases multiple runtime objects. The fourth column shows the number of static heap pointers identified by the pointer analysis, the fifth column shows the number of static global pointers identified by the pointer analysis, and the sixth column shows the number of static stack pointers identified by the pointer analysis. The last column displays the number of dangling pointers eliminated by LightDE. These data show the difference between the program’s static information and its runtime information.
5.2.2. Runtime Overhead
Figure 6. LightDE’s Runtime on SPEC CPU2006.
Figure 6 shows the runtime overhead introduced by LightDE in each project. When eliminating only heap and global dangling pointers, the geometric mean of the runtime overhead introduced by LightDE is 3.76%. When eliminating all dangling pointers, including heap, stack, and global dangling pointers, the geometric mean of the runtime overhead is 8.66%. It is worth noting that the additional time or memory overhead introduced by LightDE in some projects is far below 0.01%. To calculate the geometric mean, we set the overhead for these projects to 1% in this paper.
As shown in Figure 6, when LightDE eliminates only heap and global dangling pointers, the additional runtime overhead introduced in most projects is very low. However, we also observed significant runtime overhead in the 471.omnetpp and 483.xalancbmk projects. The reason for this is that these projects are memory allocation-intensive applications. These two programs frequently allocate and release memory, requiring frequent maintenance of the hash table and the lock-free queue (Valois, 1995), which leads to significant performance overhead. As shown in Figure 6, when LightDE eliminates stack, heap, and global dangling pointers, the additional runtime overhead increases. This is because our method requires the original thread of the program to store the memory addresses of stack pointers in our designed data structure and notify our thread to remove the corresponding local pointers upon function exit. For example, in the 453.povray and 473.astar projects, the frequent function entry and exit caused relatively high additional runtime overhead. Overall, the additional time overhead introduced by LightDE remains very low.
5.2.3. Memory Overhead
Figure 7. LightDE’s Memory Overhead on SPEC CPU2006.
Figure 8. Runtime Overhead Comparison With Dangling Pointer Elimination Methods.
Figure 7 shows the additional memory overhead introduced by LightDE in each project. The geometric mean of the memory overhead introduced by LightDE is 3.88% when eliminating only heap and global dangling pointers. When eliminating all dangling pointers, including heap, stack, and global dangling pointers, the geometric mean of the memory overhead is 7.26%.
As shown in Figure 7, when considering only heap and global pointers, the additional memory overhead introduced by LightDE is very low in most projects. However, in the projects 447.dealII, 471.omnetpp, and 482.sphinx3, LightDE introduced significantly high memory overhead. We analyzed the reasons for the high memory overhead introduced by LightDE in some projects. The memory overhead of LightDE mainly comes from two sources: the global hash table and the lock-free queue used by the program’s original thread to pass object information to the LightDE thread. LightDE iteratively eliminates dangling pointers in a separate thread. Since the results of pointer analysis can be very inaccurate under certain circumstances, it can lead to LightDE checking too many irrelevant pointers. This ineffective checking can result in excessive time consumption, causing the object information passed from the original thread to the LightDE thread to accumulate in the lock-free queue, which leads to higher memory overhead. When considering stack pointers, we observed that in the 473.astar project, a significantly high memory overhead was introduced. This is due to the frequent function calls in the 473.astar project, which generated a large number of stack pointers, resulting in very high memory consumption. Overall, the memory overhead introduced by LightDE remains very low.
5.2.4. Overhead Comparison
In this section, we compare the runtime and memory overhead of LightDE with all existing methods of eliminating dangling pointers (Lee et al., 2015; Younan, 2015; Van Der Kouwe et al., 2017; Liu et al., 2018; Shen and Dolan-Gavitt, 2020). The additional runtime overhead introduced by existing methods and LightDE is shown in Figure 8. Since FreeSentry only reports results of 7 SPEC CPU2006 projects, we display all the projects reported by FreeSentry in the left image of Figure 8. The ’?’ in the figure indicates that the corresponding overhead for the project was not reported in their method. All the data we used comes from the reports of these methods in their papers.
The geometric mean of the runtime overhead introduced by DangNull in its reported projects is 55%. It is worth noting that DangNull only considers heap pointers and does not take stack pointers and global pointers into account. FreeSentry reported the overhead of eliminating only heap dangling pointers and the overhead of eliminating both heap and stack pointers. Since FreeSentry reported too few SPEC CPU2006 projects, we do not calculate the geometric mean of its runtime overhead. The geometric mean of the runtime overhead introduced by DangSan is 41%. The geometric mean of the runtime overhead introduced by pSweeper is 17.2%. When considering only global and heap pointers, the geometric mean of the runtime overhead introduced by our method is 3.76%. When considering all pointers, the geometric mean is 8.66%. Our method introduces significantly lower runtime overhead compared to existing dangling pointer elimination methods, whether considering stack pointers or not.
When comparing memory overhead, we only compared with the better performing dangling pointer elimination methods DangSan and pSweeper. The geometric mean of the memory overhead introduced by DangSan is 240%, and the geometric mean of the memory overhead introduced by pSweeper is 247.3%. The memory overhead introduced by LightDE is significantly lower than that of existing dangling pointer elimination methods, regardless of whether stack pointers are considered.
Figure 9. LightDE’s Runtime Overhead on SPLASH-2X.
Figure 10. LightDE’s Memory Overhead on SPLASH-2X.
5.3. Scalability on Multi-threaded Applications
We used the SPLASH-2X benchmark suite to evaluate LightDE’s scalability on multithreaded applications. Figures 9 and 10 show the additional memory and runtime overhead introduced by LightDE compared to the original program. We ran all the projects in the SPLASH-2X benchmark suite. We evaluated the overhead of setting each program to run 1, 2, 4, 8, 16, 32, and 64 threads. We evaluated the overhead of using LightDE to eliminate only heap and global dangling pointers, as well as the overhead of using LightDE to eliminate stack, heap, and global dangling pointers.
Figure 9 shows the additional runtime overhead introduced by LightDE in the SPLASH-2X benchmark suite. When eliminating only heap and global dangling pointers, the geometric mean of the runtime overhead caused by LightDE is 1.23%. When eliminating all dangling pointers, including heap, stack, and global dangling pointers, the geometric mean of the runtime overhead is 8.66%. From Figure 9, it can be seen that when only heap and global dangling pointers are eliminated, LightDE introduces almost no additional runtime overhead. In this case, the runtime of the programs protected by LightDE overlaps with the runtime of the unprotected programs. However, when LightDE eliminates stack, heap, and global dangling pointers, it results in significant additional runtime overhead in the ”lu_ncb” and ”radiosity” projects. We analyzed this phenomenon. In our design, the memory addresses of stack pointers are stored in our data structure by the original program threads, which can lead to considerable time overhead. Overall, the additional runtime overhead introduced by LightDE remains very low.
Figure 10 shows the additional memory overhead introduced by LightDE, as well as the memory overhead of the original SPLASH-2X programs. When considering only global and heap pointers, the geometric mean of the memory overhead introduced by LightDE is 2.10%. When considering all pointers, including heap, stack, and global pointers, the geometric mean of the memory overhead is 2.31%. From Figure 10, it can be seen that in most of programs, the memory overhead introduced by LightDE is negligible, regardless of whether stack pointers are considered. However, the projects ”water_nsquared” and ”water_spatial” result in relatively higher memory overhead. In the ”water_nsquared” project, memory overhead increases with the number of threads, indicating that a large number of objects are created by the program’s child threads rather than by the main thread. In the ”water_spatial” project, the memory increase remains almost constant across all thread counts, suggesting that the increase in memory overhead does not correlate with the number of threads. This implies that the memory overhead primarily comes from the main thread, which creates more objects. Overall, the additional memory overhead introduced by LightDE remains very low.
6. Limitations and Future Work
We have demonstrated through experiments that LightDE is very effective at eliminating dangling pointers. However, LightDE still faces some limitations. The first limitation of LightDE comes from the imprecision of static object abstraction. As analyzed in our experiments, the reason of the high memory overhead in projects 471.omnetpp and 483.xalancbmk is the inaccuracy of the pointer analysis results. This inaccuracy is due to the fact that the object abstraction method we use in the pointer analysis is too coarse. In our pointer analysis, we abstract the memory allocation point as a single object. In a static case, when a pointer at offset $f$ of a static object $o1$ points to another static object $o2$, the corresponding runtime case is that all runtime objects created at the allocation point of $o1$ are considered to have pointers at offset $f$ that point to all runtime objects created at the allocation point of $o2$. This leads to checking too many irrelevant pointers when eliminating heap dangling pointers. This excessive checking of unrelated pointers consumes a lot of time during the elimination process, causing the LightDE thread to be unable to promptly release the objects passed by the original program thread through the lock-free queue, resulting in very high memory overhead. In our future work, we will investigate more accurate object abstraction methods to avoid excessive checks of irrelevant heap pointers during the elimination of dangling pointers.
LightDE has the issue of false positives. In some cases, programmers deliberately use dangling pointers but not for memory access. For example, some programs may use dangling pointers to calculate the address of another memory location, or convert dangling pointers to integers to compare numerical values with other integers. For instance, suppose we have two dangling pointers, $p$ and $q$, which are converted to integers and then compared using $p>q$. Since LightDE sets dangling pointers to null pointers, the comparison results change, leading to false positives. In this case, LightDE incorrectly sets the dangling pointer to null. This issue is a common problem faced by all methods that defend against UAF vulnerabilities by eliminating dangling pointers.
LightDE also has the issue of false negatives. Our method only eliminates dangling pointers in memory and cannot eliminate those in registers. When compiling a program, the compiler promotes local pointers, such as stack pointers and function parameters, to registers to improve the program’s execution speed. During program execution, register pointers reside in registers while functions are active, and when functions are inactive, register pointers are temporarily stored on the stack. Therefore, when memory is released, we cannot determine the memory space of the register pointer. Therefore, register pointers cannot be eliminated by LightDE, potentially leading to false negatives. To our knowledge, this is a common limitation of all methods that defend against UAF vulnerabilities by eliminating dangling pointers, not just ours.
7. RELATED WORK
There are two main methods for eliminating dangling pointers: The first method uses taint tracking technology to track pointers and eliminate them after the heap memory is freed. The second method involves storing pointer addresses in a specific global structure and then eliminating these pointers after the heap memory is freed. Undangle (Caballero et al., 2012) allocates a taint identifier for pointers during memory allocation and dynamically tracks the propagation of taints. When memory is freed, Undangle identifies dangling pointers through the taint identifier. This method can eliminate all dangling pointers, whether they are in memory, in registers, or have been converted to integers for propagation. However, this method incurs extremely high runtime and memory overhead, making it impractical for real-world applications and suitable only for vulnerability detection.
Another method is to store the pointer addresses in a carefully designed global structure and eliminate the corresponding dangling pointers when heap memory is freed. This method is adopted by DangNULL (Lee et al., 2015), FreeSentry (Younan, 2015), DangSan (Van Der Kouwe et al., 2017), HeapExpo (Shen and Dolan-Gavitt, 2020), and pSweeper (Liu et al., 2018). DangNULL and FreeSentry insert a specific function after pointer assignment instructions during compilation. This function is used to look up the objects pointed to by the pointers and add the pointers to the object’s pointer set. DangNULL and FreeSentry maintain the object pointed to by the pointer in real-time. When the object pointed to by a pointer changes, DangNULL and FreeSentry remove the pointer from the pointer set of the previously pointed-to object and add it to the pointer set of the newly pointed-to object. DangNULL uses a tree structure to store the object’s pointer set, while FreeSentry uses a two-level mapping table. DangSan uses a system log-like method to record the objects pointed to by pointers. DangSan inserts a function after pointer assignment instructions to add the pointer to the pointer set of the object it points to. If the target of a pointer changes, it does not remove the pointer from the set of the previously pointed-to object. This approach avoids the real-time maintenance of object pointer sets, resulting in lower performance overhead. When an object is freed, DangSan checks whether the pointers in the object’s pointer set still point to the freed object. HeapExpo uses the same method as DangSan but prohibits the compiler from optimizing local pointers into registers, resulting in very high performance overhead. To further improve performance, pSweeper adopts an asynchronous method to eliminate dangling pointers. During program execution, pSweeper only stores pointers in its carefully designed global data structure, avoiding the need to look up the objects pointed to by the pointers. pSweeper runs a separate thread on an idle CPU core. This thread continuously scans the pointers recorded in the global data structure to check if they point to freed memory and eliminates dangling pointers. However, all these methods require interrupting the program’s execution flow to store the memory addresses of pointers in their global data structures and to look up the objects pointed to by the pointers.
8. Conclusion
This paper proposes LightDE, a lightweight method for eliminating dangling pointers to defend against UAF vulnerabilities. LightDE stores pointer points-to information in the program’s data segment at compile time based on our pointer analysis results, avoiding the need to store pointer memory addresses and look up pointer targets during program execution. To achieve this, we present a structure-sensitive pointer analysis method for analyzing the type information of objects. Additionally, we introduce a method for mapping static pointer analysis results to runtime information to accurately eliminate dangling pointers. We implemented a prototype system of LightDE and validated its effectiveness and efficiency.